# Probing the Origins I: Generalised Additive Model inference of birth radii for Milky Way stars in the solar vicinitythanks: Catalogue reference

### v1

@@ 5 Data availability
The catalogue is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Appendix A Model using the mgcv package
The effective degrees of freedom (EDF), however, are controlled during fitting by methods such as GCV (Generalised Cross-Validation), AIC (Akaike Information Criterion), or REML (Restricted Maximum Likelihood, as previously described and used in the current paper). This means that the exact choice of $k$ is not overly critical as long as it is sufficiently large to capture the underlying data pattern while remaining computationally efficient. As noted in the documentation, choosing a higher $k$ value can help prevent underfitting, but excessively high values may result in overfitting or unnecessary computational cost.
@@
The scatterplot of the response variable versus the fitted values from the GAM model in Fig. 22 (generated through the gam.check command in the mgcv package) demonstrates the model’s overall effectiveness in reproducing observed data. Across most radii, the points align closely along the 45-degree line, indicating a strong agreement between the fitted and observed values. The consistent spread of points suggests homoscedasticity, further validating the model’s adequacy. However, minor deviations are evident at very small radii (below 2.5 kpc), where the model exhibits degeneracy. This behaviour likely stems from the limited resolution of the original chemical enrichment models and the sensitivity of the logarithmic link function (link=’log’) to values approaching zero, which can introduce instability. Despite these limitations, the GAM performs exceptionally well across the broader range of $t$ and [Fe/H], reinforcing its robustness and supporting the conclusion that these stars were likely formed at even smaller radii. Far from compromising our findings, this slight overestimation of $R_{\rm b}$ at near-zero radii strengthens our key result: stars have migrated from the innermost regions of the Galaxy to the solar neighbourhood. Figure 22 also provides insight into these minor discrepancies observed in Fig. 5, tracing their origin to resolution limitations.
@@
Appendix B Comparison between GAM models for [Fe/H] and [Mg/H]
In this appendix, we present a supplementary analysis showing the results from the GAM regression using [Mg/H] and compare them with the results obtained using [Fe/H]. We chose to run the GAM with [Mg/H] instead of [Mg/Fe] to fully remove the influence of Fe in this auxiliary model. Figures 23 and 24 are analogous to Figs. 5 and 6 in the main text, but use [Mg/H] as an independent variable in the GAM model instead of [Fe/H]. Figures 25 and 26 compare the $\langle R_{\rm{b}}\rangle$ estimates obtained using the main GAM (i.e., with [Fe/H] as an independent variable) and the alternative GAM using [Mg/H].
@@ Appendix B Comparison between GAM models for [Fe/H] and [Mg/H]
Figure 23: Similar to Fig. 5, but with a model considering [Mg/H], instead of [Fe/H].
Figure 24: Similar as Fig. 6, but instead of depicting [Fe/H] in the $x$-axis, depicting [Mg/H].
@@ C.2 Potential blurred thin disc stars classified as halo stars
Figure 29: The $\langle Z_{\rm{max}}\rangle$ and $\langle e\rangle$ of the stars removed from the main analysis using the Toomre diagram are shown here. Stars classified as most likely belonging to the Galactic halo are coloured magenta, consistent with previous figures, while stars coloured according to the sample groups represent potential blurred disc stars. The marker size is proportional to $|\Delta R|$. The adjacent plots display the distributions of $\langle Z_{\rm{max}}\rangle$ and $\langle e\rangle$, with median values indicated by a dotted line for potential halo stars, a solid line for potential disc stars, and the dashed line for stars with low $W$ or $U$. As in Fig. 15, the scale heights for the thin and thick discs are shown as dotted and dot-dashed lines, respectively (i.e. 300 and 900 pc; McMillan 2017). Note that while the $x$-axis is the same across all subplots, the $y$-axis scales differ to better depict the distribution of each subplot individually. Unlike similar figures, we omit the annotated medians in the adjacent distribution plots to avoid information overload.
@@ C.3 Heliocentric distances
C.4 A few different perspectives on $\langle R_{\rm{b}}\rangle$ vs. $\langle R_{\rm{g}}\rangle$, $\langle R_{\rm{apo}}\rangle$, and $\langle R_{\rm{peri}}\rangle$
@@
Figure 31: Similar to Fig. 12, but contrasting $\langle R_{\rm{peri}}\rangle$ instead of $\langle R_{\rm{g}}\rangle$.
Figure 32: Similar to Fig. 12, but contrasting $\langle R_{\rm{apo}}\rangle$ instead of $\langle R_{\rm{g}}\rangle$. In this case, we changed the range of $\langle R_{\rm{apo}}\rangle$, to better depict all the stars.
Figure 33: Similar to Fig. 12, but with the movement classification. Stars with unchanged orbital radii are marked in grey, whereas those consistent with churning keep their colour scheme used throughout this paper.
Figure 34: Scatterplot of $\langle R_{\rm{b}}\rangle$ vs. stellar age ($\overline{t}_{\star}$) for all groups in our sample. A gray horizontal dashed line at 3 kpc highlights the approximate $\langle R_{\rm{b}}\rangle$ limit for stars that are not super-metal-rich. Group 2, consisting of super-metal-rich stars, is notably not constrained by this limitation.
@@ C.5 Angular momentum in Cartesian and cylindrical coordinates
Figure 36: Distribution of $L_{x}$ for the stars in our sample. Each panel represents one of the HC metallicity-stratified groups, arranged in order of decreasing $\langle\rm{[Fe/H]}\rangle$. Within each panel, the $L_{x}$ distributions for blurred/undisturbed stars (marked as “Equal”), outward-migrating stars, and inward-migrating stars are shown in orange, blue, and green, respectively. Gaussian kernel density estimates are overlaid to smooth the binning effects. Vertical dashed lines indicate the median values for each distribution in the corresponding colours.
@@
Figure 39: Same as Fig. 18, but in Cartesian coordinates.

### v2

@@ 5 Data availability
The catalogue is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/696/A205. Additional parameters can be provided upon reasonable request; please contact the corresponding author for such inquiries.
@@ Appendix A Model using the mgcv package
The effective degrees of freedom, however, are controlled during fitting by methods such as generalised cross-validation (GCV), akaike information criterion (AIC), or restricted maximum likelihood (REML, as previously described and used in the current paper). This means that the exact choice of $k$ is not overly critical as long as it is sufficiently large to capture the underlying data pattern while remaining computationally efficient. As noted in the documentation, choosing a higher $k$ value can help prevent underfitting, but excessively high values may result in overfitting or unnecessary computational cost.
@@
The scatterplot of the response variable versus the fitted values from the GAM in Fig. 22 (generated through the gam.check command in the mgcv package) demonstrates the model’s overall effectiveness in reproducing observed data. Across most radii, the points align closely along the 45-degree line, indicating a strong agreement between the fitted and observed values. The consistent spread of points suggests homoscedasticity, further validating the model’s adequacy. However, minor deviations are evident at very small radii (below 2.5 kpc), where the model exhibits degeneracy. This behaviour likely stems from the limited resolution of the original chemical enrichment models and the sensitivity of the logarithmic link function (link=‘log’) to values approaching zero, which can introduce instability. Despite these limitations, the GAM performs exceptionally well across the broader range of $t$ and [Fe/H], reinforcing its robustness and supporting the conclusion that these stars were likely formed at even smaller radii. Far from compromising our findings, this slight overestimation of $R_{\rm b}$ at near-zero radii strengthens our key result: stars have migrated from the innermost regions of the Galaxy to the solar neighbourhood. Figure 22 also provides insight into these minor discrepancies observed in Fig. 5, tracing their origin to resolution limitations.
@@
Appendix B Comparison between generalised additive models for [Fe/H] and [Mg/H]
In this appendix, we present a supplementary analysis showing the results from the GAM regression using [Mg/H] and compare them with the results obtained using [Fe/H]. We chose to run the GAM with [Mg/H] instead of [Mg/Fe] to fully remove the influence of Fe in this auxiliary model. Figures 23 and 24 are analogous to Figs. 5 and 6 in the main text, but use [Mg/H] as an independent variable in the GAM instead of [Fe/H]. Figures 25 and 26 compare the $\langle R_{\rm{b}}\rangle$ estimates obtained using the main GAM (i.e. with [Fe/H] as an independent variable) and the alternative GAM using [Mg/H].
@@ Appendix B Comparison between generalised additive models for [Fe/H] and [Mg/H]
Figure 23: Similar to Fig. 5 but with a model considering [Mg/H], instead of [Fe/H].
Figure 24: Similar to Fig. 6, but instead of depicting [Fe/H] in the $x$-axis, it depicts [Mg/H].
@@ C.2 Potential blurred thin disc stars classified as halo stars
Figure 29: The $\langle Z_{\rm{max}}\rangle$ and $\langle e\rangle$ of the stars removed from the main analysis using the Toomre diagram are shown here. Stars classified as most likely belonging to the Galactic halo are coloured magenta, consistent with previous figures, while stars coloured according to the sample groups represent potential blurred disc stars. The marker size is proportional to $|\Delta R|$. The adjacent plots display the distributions of $\langle Z_{\rm{max}}\rangle$ and $\langle e\rangle$, with median values indicated by a dotted line for potential halo stars, a solid line for potential disc stars, and the dashed line for stars with low $W$ or $U$. As in Fig. 15, the scale heights for the thin and thick discs are shown as dotted and dot-dashed lines, respectively (i.e. 300 and 900 pc; McMillan 2017). We note that while the $x$-axis is the same across all subplots, the $y$-axis scales differ to better depict the distribution of each subplot individually. Unlike similar figures, we omit the annotated medians in the adjacent distribution plots to avoid information overload.
@@ C.3 Heliocentric distances
C.4 A few different perspectives on $\langle R_{\rm{b}}\rangle$ versus $\langle R_{\rm{g}}\rangle$, $\langle R_{\rm{apo}}\rangle$, and $\langle R_{\rm{peri}}\rangle$
@@
Figure 31: Similar to Fig. 12 but contrasting $\langle R_{\rm{peri}}\rangle$ instead of $\langle R_{\rm{g}}\rangle$.
Figure 32: Similar to Fig. 12 but contrasting $\langle R_{\rm{apo}}\rangle$ instead of $\langle R_{\rm{g}}\rangle$. In this case, we changed the range of $\langle R_{\rm{apo}}\rangle$, to better depict all the stars.
Figure 33: Similar to Fig. 12 but with the movement classification. Stars with unchanged orbital radii are marked in grey, whereas those consistent with churning keep their colour scheme used throughout this paper.
Figure 34: Scatterplot of $\langle R_{\rm{b}}\rangle$ versus stellar age ($\overline{t}_{\star}$) for all groups in our sample. A grey horizontal dashed line at 3 kpc highlights the approximate $\langle R_{\rm{b}}\rangle$ limit for stars that are not super-metal-rich. Group 2, consisting of super-metal-rich stars, is notably not constrained by this limitation.
@@ C.5 Angular momentum in Cartesian and cylindrical coordinates
Figure 36: Distribution of $L_{x}$ for the stars in our sample. Each panel represents one of the HC metallicity-stratified groups, arranged in order of decreasing $\langle\rm{[Fe/H]}\rangle$. Within each panel, the $L_{x}$ distributions for blurred/undisturbed stars (marked as ‘Equal’), outward-migrating stars, and inward-migrating stars are shown in orange, blue, and green, respectively. Gaussian kernel density estimates are overlaid to smooth the binning effects. Vertical dashed lines indicate the median values for each distribution in the corresponding colours.
@@
Figure 39: Same as Fig. 18 but in Cartesian coordinates.